The probability that an international flight leaving the United States is delayed in departing (event D) is .29. The probability that an international flight leaving the United States is a transpacific flight (event P) is .59. The probability that an international flight leaving the U.S. is a transpacific flight and is delayed in departing is .11. (a) What is the probability that an international flight leaving the United States is delayed in departing given that the flight is a transpacific flight? (Round your answer to 4 decimal places.) Probability (b) In this problem, are D and P independent? Yes No
Answer #1

from above P(D)=0.29 , P(P) =0.59 , P(D n P)=0.11

a)probability that an international flight leaving the United States is delayed in departing given that the flight is a transpacific flight =P(D|P) =P(D n P)/P(P) =0.11/0.59 =0.1864

b)

No, since P(D|P) is not equal to P(D)